OpenStudy (anonymous):

There are many ropes keeping a hot air balloon from floating away before a balloon race. One of these ropes is fixed to the ground at a 45° angle. Another is fixed to the ground at a 30° angle. If the hot air balloon is 18 feet off the ground, what is the distance between the ground directly underneath the balloon and the second rope rounded to the nearest hundredth of a foot? posting pic one sec plzz help i thing i got it but not sure

OpenStudy (amistre64):

a tangent application would help out alot

OpenStudy (amistre64):

assuming level ground that is

OpenStudy (amistre64):

also, knowing the basic trig triangles would amount to the same thing i believe

OpenStudy (amistre64):

| | 1 t=30 | ---------- sqrt(3) since everything 1 is sacled by a factor of 18, all the linear stuff is scaled by 18 all the same

OpenStudy (anonymous):

31.18 tan 30=18/x x=18/tan30